OpenStudy (anonymous):

Find dy/dx by implicit differentiation. 4+x=sin(xy2)

OpenStudy (anonymous):

I got to 1-cos(y^2)=cos(x2yy') I'm not sure what to do next

OpenStudy (baldymcgee6):

[4+x]' = [sin(2xy)]' 1 = cos(2xy) *[2xy]'